# DirectMultiStep: Direct Route Generation for Multi-Step Retrosynthesis

### v1

@@ Abstract
Traditional computer-aided synthesis planning (CASP) methods rely on iterative single-step predictions, leading to exponential search space growth that limits efficiency and scalability. We introduce a transformer-based model that directly generates multi-step synthetic routes as a single string by conditionally predicting each molecule based on all preceding ones. The model accommodates specific conditions such as the desired number of steps and starting materials, outperforming state-of-the-art methods on the PaRoutes dataset with a 2.2x improvement in Top-1 accuracy on the n1 test set and a 3.3x improvement on the n5 test set. It also successfully predicts routes for FDA-approved drugs not included in the training data, showcasing its generalization capabilities. While the current suboptimal diversity of the training set may impact performance on less common reaction types, our approach presents a promising direction towards fully automated retrosynthetic planning.
@@ 1 Introduction
A significant contribution to the development of algorithmic frameworks for identifying synthetic routes leading to any desired compound (subject of the 1990 Nobel Prize in Chemistry) was made by Elias James Corey. Corey’s framework, now known as retrosynthetic analysis, begins with identifying atoms that would serve as reaction centers. Disconnecting bonds between these centers results in the formation of hypothetical fragments (called synthons) from which a precursor molecule can be created. This mapping from fragments to actual molecules is one-to-many because there is usually more than one functional group that could participate in any given type of reaction. Importantly, meticulous application of Corey’s framework (i.e., systematically breaking small subsets of bonds) will eventually lead to commercially available starting materials. The algorithmic nature of this process allowed Corey to envision automating these rules to create Computer-Aided Synthesis Planning (CASP) as early as 1969 [1].
@@
Recent advancements in data science and machine learning (ML) methods have led to a surge of interest in developing CASP methods [2, 3]. The vast majority of existing methods [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35] are designed to automate single-step retrosynthetic (SSR) analysis, i.e. predicting a list of compounds from which a target product could be made in one step. A full multi-step route can be determined by iteratively applying SSR methods to the identified precursors until a termination condition is satisfied (e.g., identifying reactants in the database of commercially available compounds). Notably, because each SSR method call creates a list of candidates, iterative application of these methods generates an exponentially growing search space. Graph traversal algorithms such as Monte Carlo Tree Search (MCTS), Depth-First Proof Number (DFPN) search, and A-star (Retro*) have been applied to efficiently traverse this exponential search space [36, 37, 38].
@@
However, evaluation of performance of these SSR methods on full route prediction has been limited by the scarcity of open-source datasets containing valid multi-step routes. The creation of PaRoutes [39], an open-source dataset containing 450k multi-step routes extracted from the United States Patent and Trademark Office (USPTO), marked a major development. Notably, even state-of-the-art SSR models [6, 40] combined with advanced search algorithms (MCTS, DFPN, Retro*) correctly identify multi-step routes (Top-1 accuracy) for only 17% and 10% of the target compounds in n1 and n5, respectively. This performance can be rationalized by recognizing that retrosynthesis is inherently a multi-step problem: the optimal choice of reaction to make a compound depends on subsequent steps in the synthesis. For example, a common pattern in multi-step routes includes: 1) protection of certain functional groups, 2) the desired transformation of unprotected groups, and 3) deprotection of the protected groups. An SSR method applied to the reactant of reaction 2 may output numerous candidate precursors; however, knowing that protective groups are removed in subsequent steps dramatically changes the probability distribution over those precursors.
In this work, we propose a novel approach for direct prediction of multi-step routes, bypassing the need for single-step models and sophisticated exponential graph traversal algorithms. Our DirectMultiStep model demonstrates state-of-the-art performance on n1 and n5 evaluation sets and shows generalizability by successfully predicting routes for FDA-approved drugs absent from the training set. We also discuss the model’s limitations, which may include the need to split longer routes into sub-routes, provide the structure of the starting material, or supply a precursor to the target compound when dealing with transformations not well-represented in the training dataset. Despite these limitations, we believe our results are promising enough to warrant further attention to developing retrosynthetic methods that directly predict multi-step routes.
2 Methods
2.1 Definitions and multi-step route representation
@@
This subsection defines key terms related to our multi-step retrosynthetic methodology. The "target compound" (green in Fig. 1) represents the final product of the multi-step synthesis tree. Starting materials (SM, orange in Fig. 1) are compounds for which no further precursors need to be identified, i.e., they are the leaves of the synthesis tree. The "number of steps" denotes the largest number of reactions from SM to the target compound (tree height). In our approach, routes are represented as recursive dictionaries (Fig. 1) containing the SMILES representation of the molecule and a list of other dictionaries, containing either starting materials or trees leading to precursors of the current node. Removing space and newline characters creates a string representation predicted by our transformer model. We train the DirectMultiStep (DMS) transformer, which takes the target compound, (optionally) starting material, and number of steps as input, and predicts the multi-step route represented as a string.
2.2 Data curation and preparation
A second version of the PaRoutes [39] dataset containing 450k routes is stripped of all metadata and stored as recursive dictionaries representing multi-step routes. The two evaluation sets, n1 and n5, are processed similarly. A training partition is created by removing all permutations (swapping left and right subtrees) of routes in n1 and n5 from the full dataset, resulting in 407k routes. The training dataset is augmented by adding 2 permutations for each route. To train the DMS model which takes starting material as an input, we find all starting materials (leaves) for each tree and store a combination of the target compound with each starting material as a separate entry. As a result, DMS with SM is trained on 3 349 118 inputs, and the DMS without SM on 1 078 951 inputs. All SMILES strings are tokenized by treating each character as a token, and the string representation of the multi-step route is tokenized similarly, treating delimiters of the tree (smiles’, children’, [’, ]’, {’, and }’) as separate tokens. The final vocabulary size is 52 (including start, end, and padding tokens), the largest multi-step route has 1074 tokens, while the largest target compound and starting material have 145 and 135 tokens, respectively.
@@ 2.2 Data curation and preparation
Figure 1: The workflow of DirectMultiStep. (a) The SMILES representation of the target compound (green), (optional) starting material (orange), and the number of steps are tokenized, concatenated, and fed into our transformer model. The model predicts a string representation of the multi-step synthesis tree. Spaces are added for clarity, and indentations indicate the levels in the synthesis route (tree). (b) Molecular structures corresponding to the target compound (green), (optional) starting material (orange), and the predicted synthesis tree with structures of all molecules.
Figure 2: Distribution of the relative frequencies of route lengths (in terms of number of steps) in the training dataset before augmentation with permutations (407 148 routes, black), n1 test set (10 000 routes, blue), and n5 test set (10 000 routes, purple). Distribution is split into routes shorter (left subplot) and longer than 6 steps (right subplot).
2.3 Model architecture and training
We present two variants of the DirectMultiStep (DMS) model: DMS-10M and DMS-60M, with approximately 10 million and 60 million parameters, respectively. Both models employ the transformer encoder-decoder architecture [41]. DMS-10M has 6 layers and 8 attention heads in both encoder and decoder, with inputs embedded in 256 dimensions and amplified 3x in the feedforward blocks. DMS-60M has 8 layers and 8 attention heads, with inputs embedded in 512 dimensions and amplified 4x in the feedforward blocks. Both models use GeLU activations [42]. The encoder of DMS with SM receives a concatenation of the tokenized target compound, starting material, and number of steps, while the encoder of DMS without SM receives only the tokenized target compound and number of steps. In all models, the decoder receives output from the encoder and the start token of the multi-step route.
During training, a 10% dropout rate is used in embedding, attention, and feedforward layers. The ADAM optimizer [43] is used with a learning rate scheduler that warms up from 0 to $3\times 10^{-4}$ for DMS-10M and $1\times 10^{-4}$ for DMS-60M during the first 10% of training steps and undergoes cosine decay to $3\times 10^{-5}$ over 20 epochs. Gradients are clipped to 1.0. Finally, during training, each token in the input to the encoders has a 5% chance of being masked. Training requires approximately 200 clock hours using 4 NVIDIA A100-40 GB GPUs on an external cluster. An additional 450 hours were used for experimentation with model architectures and optimization of hyperparameters.
2.4 Beam search for route generation
We utilize beam search with a width of 50 to generate several candidates for the same target compound (and starting material, if provided). The cumulative probability of each beam serves as the ranking metric for prediction accuracy (Top-$K$). After beam search, predicted routes are checked for the validity of all SMILES strings (which are canonicalized), the correctness of the target compound, the presence of starting materials in the stock compound (if applicable), and the absence of repetitions.
@@ 2.5 Evaluation metrics and post-processing
Evaluation is performed on all 10 000 routes in both n1 and n5, with the deepest leaf selected as the starting material to be included in the input to the encoder (for models that use starting materials). Top-$K$ accuracy evaluates the percentage of top $K$ predictions precisely matching correct multi-step routes from the test set. Predictions matching any permutation of the correct route are counted as matches.
@@ 3 Results and Discussion
3.1 Test accuracy
Top-$K$ accuracy is shown in Table 1 for set-n1 and Table 2 for set-n5. Our DirectMultiStep-SM-10M model achieves state-of-the-art performance, surpassing all previous methods in all Top-$K$ accuracies, with the most significant increase in Top-1 accuracy: 2.2x on n1 and 3.3x on n5. Remarkably, our model’s Top-3 accuracy exceeds the Top-10 accuracy of top tree search methods using single-step retrosynthesis (SSR), highlighting the advantages of the multi-step first approach.
The DirectMultiStep-SM-60M model, despite having six times more parameters than the 10M variant, exhibits lower performance on higher Top-$K$ accuracies, suggesting that the 10M model size is optimal for the current dataset when starting material structure is provided. However, achieving comparable performance without starting material information requires a larger model, as demonstrated by the DirectMultiStep-noSM-60M model’s performance. These findings reveal a trade-off between model generality and size at the current dataset scale: less input information necessitates a larger model to maintain high performance. In all subsequent analysis, we use the routes predicted by the DirectMultiStep-SM-10M model.
@@ 3.1 Test accuracy
DirectMultiStep-SM-10M 0.38 0.46 0.50 0.52 0.53 0.56
DirectMultiStep-SM-60M 0.38 0.46 0.48 0.50 0.51 0.52
DirectMultiStep-noSM-60M 0.36 0.42 0.44 0.45 0.45 0.46
@@
DirectMultiStep-SM-10M 0.33 0.40 0.43 0.44 0.46 0.48
DirectMultiStep-SM-60M 0.34 0.41 0.43 0.44 0.45 0.46
DirectMultiStep-noSM-60M 0.32 0.37 0.39 0.40 0.41 0.41
aThese data are collected from the 2.0 version of PaRoutes in their GitHub repository [39]
Importantly, our model requires only one call to obtain multiple routes for each target compound, instead of hundreds of single-step model calls and even more reaction template applications (as shown in PaRoutes [39]). Furthermore, our model’s inference time (beam size 50) on a single GPU is comparable (3-16 seconds depending on the number of steps) to the time required to find the first successful route using previous methods (7-50 seconds [39])
Fig. 3 shows the distribution of Top-1 and Top-10 accuracy over different route lengths on both n1 and n5. Given that 90% of routes in the training partition have 4 or fewer steps Fig. 2), one would expect accuracy to decrease dramatically with increasing route length. However, the performance on routes with 5-8 steps is comparable to that of shorter routes, and the performance on 9-step routes is even comparable to that of 2-step routes. The exceptional accuracy on 9-step routes could be rationalized by the fact that their relative abundance in n1 and n5 does not exceed the relative abundance in the training partition (unlike for routes with length 3-7, as seen in Fig. 3).
@@
Figure 3: Distribution of Top-1 and Top-10 accuracy of predictions with our model on test sets n1 and n5. There is only one route with length 10 in n1, and our model does not predict it correctly. That route is reproduced by splitting it in half, as shown in Fig. 4

### v2

# DirectMultiStep: Direct Route Generation for Multistep Retrosynthesis

@@ Abstract
Traditional computer-aided synthesis planning (CASP) methods rely on iterative single-step predictions, leading to exponential search space growth that limits efficiency and scalability. We introduce a series of transformer-based models, utilizing mixture of experts approach, that directly generate multistep synthetic routes as a single string by conditionally predicting each molecule based on all preceding ones. Our models can accommodate specific conditions such as the desired number of steps and starting materials, with the top-performing DMS-Flex (Duo) surpassing state-of-the-art methods on the PaRoutes dataset with a 2.5x improvement in Top-1 accuracy on the n1 test set and a 3.9x improvement on the n5 test set. It also successfully predicts routes for FDA-approved drugs not included in the training data, showcasing its generalization capabilities. While the current suboptimal diversity of the training set may impact performance on less common reaction types, our multistep-first approach presents a promising direction towards fully automated retrosynthetic planning.
@@ 1 Introduction
Algorithmic frameworks for identifying synthetic routes were envisioned by Vleduts [1] and further formalized by Elias James Corey (subject of the 1990 Nobel Prize in Chemistry) into what is now known as retrosynthetic analysis. This framework begins with identifying atoms that would serve as reaction centers. Disconnecting bonds between these centers results in the formation of hypothetical fragments (called synthons) from which a precursor molecule can be created. This mapping from synthons to actual molecules is one-to-many because there is usually more than one functional group that could participate in any given type of reaction. Importantly, meticulous application of Corey’s framework (i.e., systematically breaking small subsets of bonds) will eventually lead to commercially available starting materials. The algorithmic nature of this process allowed Corey to envision automating these rules to create Computer-Aided Synthesis Planning (CASP) as early as 1969 [2].
@@
Recent advancements in data science and machine learning (ML) methods have led to a surge of interest in developing CASP methods [3, 4]. The vast majority of existing methods [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36] are designed to automate single-step retrosynthetic (SSR) analysis, i.e. predicting a list of compounds from which a target product could be made in one step. A full multistep route can be determined by iteratively applying SSR methods to the identified precursors until a termination condition is satisfied (e.g., identifying reactants in the database of commercially available compounds). Notably, because each SSR method call creates a list of candidates, iterative application of these methods generates an exponentially growing search space. Graph traversal algorithms such as Monte Carlo Tree Search (MCTS), Depth-First Proof Number (DFPN) search, and A-star (Retro*) have been applied to efficiently traverse this exponential search space [37, 38, 39]. More sophisticated search designs include utilization of hyper-graph [40], graph-based neural networks [41], reinforcement learning [42], or incorporation of previous reactions as context for single-step predictions [43].
@@
Evaluation of performance of these SSR methods on full route prediction has been limited by the scarcity of open-source datasets containing valid multistep routes. The creation of PaRoutes [44], an open-source dataset containing 450k multistep routes (163k excluding duplicates and permutations, along with two test sets of 10k routes each) extracted from the United States Patent and Trademark Office (USPTO), marked a major development. Notably, even state-of-the-art SSR models [7, 45] combined with advanced search algorithms (MCTS, DFPN, Retro*) correctly identify multistep routes (Top-1 accuracy) for only 17% and 10% of the target compounds in n1 and n5 test sets, respectively. Such performance could be rationalized by recognizing that retrosynthesis is inherently a multistep problem: the optimal choice of reaction to make a compound depends on subsequent steps in the synthesis. For example, a common pattern in multistep routes includes: (1) protection of certain functional groups, (2) the desired transformation of unprotected groups, and (3) deprotection of the protected groups. An SSR method applied to the reactant of reaction (2) may output numerous candidate precursors; however, knowing that protective groups are removed in subsequent steps dramatically changes the probability distribution over those precursors.
In this work, we propose a novel approach for direct prediction of multistep routes, bypassing the need for single-step models and sophisticated exponential graph traversal algorithms. We begin with a brief overview of the methodology and evaluation metrics. We proceed by demonstrating and discussing performance of DMS models on diverse n1 and n5 evaluation sets (10 000 routes each). We demonstrate further generalizability by successfully predicting experimental synthetic routes for FDA-approved drugs. We conclude by discussing the limitations of our approach. A more elaborate description of model architectures and rationale for certain design choices is given at the end of the paper.
2 Methodology Overview
2.1 Definitions
Target compound (blue in Fig. 1) is the final (desired) product of the multistep synthesis tree. Starting material, SM (red in Fig. 1) is a compound for which no further precursors need to be identified, i.e., a leaf of the synthesis tree. The number of steps is the largest number of reactions from SM to the target compound (tree height).
2.2 Route Representation
@@
Routes are represented as recursive dictionaries (Fig. 1) containing the SMILES representation of the molecule and a list of other dictionaries, containing either starting materials or trees leading to precursors of the current node. Removing space and newline characters creates a string representation. Our models predict that string by taking SMILES of the target compound, (optionally) starting material, and (optionally) number of steps as input.
@@
Figure 1: The workflow of DirectMultiStep. (a) The SMILES representation of the target compound (blue), (optional) starting material (red), and the number of steps are tokenized, concatenated, and fed into our transformer model. The model predicts a string representation of the multistep synthesis tree. Spaces are added for clarity, and indentations indicate the levels in the synthesis route (tree). (b) Molecular structures corresponding to the target compound (blue), (optional) starting material (red), and the predicted synthesis tree with structures of all molecules.
@@ 2.3 Evaluation Metrics and Post-Processing
Evaluation is performed on n1 and n5 (10 000 routes each) evaluation sets from PaRoutes [44], which have been created by selecting the most diverse routes from the USPTO. Fig. 2 shows the distribution of route lengths in training set, n1, and n5. The distribution in both n1 and n5 is shifted towards longer routes; for example, the fraction of routes with 4 steps or more is 15% for training set, 27% for n1 and 51% for n5.
Figure 2: Distribution of the relative frequencies of route lengths (in terms of number of steps) in the training dataset before augmentation with permutations (163 689 routes, black), n1 test set (10 000 routes, blue), and n5 test set (10 000 routes, purple). Distribution is split into routes shorter (left subplot) and longer than 6 steps (right subplot).
The structure of target compound, starting material (optionally, the deepest leaf of the correct route is chosen during evaluation), and number of steps is passed to the transformer encoder. A beam search with a width of 50 is employed to generate 50 candidate routes by sampling next tokens from the decoder. Within beam search, during the prediction of each token the cumulative log likelihoods are normalized by the sequence length at the time the new token is sampled. This adjustment ensures that shorter routes with simpler compounds are not disproportionately favored while maintaining a balance between route length and overall sequence probability. After beam search finds end-of-sequence tokens for all 50 candidates, predicted routes are checked for validity of all SMILES strings (which are canonicalized), the presence of starting materials in the stock set (following procedure of PaRoutes taken to be the set of all leaves in n1 or n5, and absence of repetitions (which can be present as permutations of other routes). Remaining routes are compared against the true (experimental) route. The reproduction of experimental routes is presented as Top-$K$ accuracy, i.e. presence of correct route ranked $K$ or lower. Importantly, the Top-$K$ accuracy is evaluated against all 10 000 targets, not just the ones for which routes are found.
2.4 Models Presented
As a baseline, we train a conventional transformer-based sequence-to-sequence model with 6 encoder and 6 decoder layers (DMS-Flash) with structures of starting materials provided during training. To achieve comparable performance without starting material as a hint, a significantly larger model with 12 encoder and 36 decoder layers is needed (DMS-Deep). We also present the first, to our knowledge, application of mixture-of-experts (MoE [46, 47, 48]) architecture to a chemical problem. To create MoE model, one replaces the 2-layer MLP that follows every attention block with an ensemble of 2-layer MLPs (called experts). A preceding router determines (for each input) the $n$-most relevant (active) experts, and passes the input through each of them. Output of each active expert is added with weights determined by the router. Notably, the number of experts active during inference does not have to be equal to the number of experts active during training, although the latter tends to be optimal. We present two MoE models, the DMS-Flex (trained with SM) and DMS-Wide (trained without SM). The relevant architectural details of models are given in Table 1.
Table 1: Specifications of architecture for different model variants.
DMS Model Encoder Decoder Hidden FF Total Active Experts
Variant Layers Layers Dimension Multipliera Experts Training Generation
Flash 6 6 256 3 1 1 1
Explorer 6 6 256 3 3 2 2
Explorer XL 6 24 256 3 3 2 2
Flex (Mono) 6 6 256 3 3 2 1
Flex (Duo) 6 6 256 3 3 2 2
Deep 12 36 256 3 1 1 1
Wide 12 12 256 3 3 2 2
a an integer by which the hidden dimension is multiplied in feed-forward layers after attention.
@@ 3 Results and Discussion
3.1 Top-K Accuracy on Evaluation Sets
3.1.1 Vanilla Transformers
Top-$K$ accuracy is shown in Table 2 for set-n1 and Table 3 for set-n5. The smallest DMS-Flash model (which receives structure of SM during generation) achieves state-of-the-art performance, surpassing all previous methods in all Top-$K$ accuracies, with the most significant increase in Top-1 accuracy: 2.4x on n1 and 3.7x on n5. To achieve comparable performance without providing the structure of the SM, a significantly larger model DMS-Deep is needed. This establishes a trade-off between size of the model (and thus generation speed) and the amount of information it receives. The limitations of the necessity to provide the structure of the SM are discussed in Section 3.1.3. Quite remarkably, a Flash model trained with structures of SM may generate routes even when SMs are not provided (Tables 2-3). While that results in a 20% drop in Top-1 accuracy, it is significantly better than a model with the same architecture that is trained without SM (Tables S6-S7). In other words, providing structure of a correct starting material helps the model learn better and improves performance even if the starting material is not used during evaluation.
@@ 3.1.1 Vanilla Transformers
DMS SM SM
Models Size Steps (Train) (Gen.) Top-1 Top-2 Top-3 Top-4 Top-5 Top-10
Flash 10M 0.41 0.50 0.54 0.56 0.58 0.60
Flash 10M 0.32 0.39 0.42 0.44 0.45 0.47
Explorer 19M 0.36 0.45 0.48 0.51 0.52 0.55
Explorer 19M 0.27 0.33 0.36 0.37 0.39 0.40
Explorer-XL 50M 0.32 0.40 0.44 0.45 0.46 0.48
Flex (Mono) 19M 0.32 0.40 0.43 0.44 0.45 0.48
Flex (Duo) 19M 0.43 0.52 0.55 0.57 0.58 0.61
Flex (Duo) 19M 0.34 0.41 0.44 0.45 0.46 0.47
Deep 41M 0.36 0.44 0.47 0.49 0.50 0.52
Wide 38M 0.38 0.46 0.49 0.50 0.51 0.53
@@
DMS SM SM
Models Size Steps (Train) (Gen.) Top-1 Top-2 Top-3 Top-4 Top-5 Top-10
Flash 10M 0.37 0.44 0.47 0.49 0.51 0.53
Flash 10M 0.30 0.36 0.39 0.40 0.41 0.43
Explorer 19M 0.33 0.40 0.43 0.45 0.46 0.49
Explorer 19M 0.25 0.31 0.34 0.35 0.36 0.37
Explorer-XL 50M 0.31 0.38 0.41 0.42 0.43 0.45
Flex (Mono) 19M 0.28 0.34 0.36 0.38 0.39 0.41
Flex (Duo) 19M 0.39 0.46 0.49 0.51 0.52 0.54
Flex (Duo) 19M 0.32 0.38 0.40 0.42 0.42 0.44
Deep 41M 0.33 0.40 0.43 0.45 0.46 0.48
Wide 38M 0.35 0.42 0.44 0.46 0.47 0.49
Remarkably, the Top-2 accuracy of DMS-Flash and the Top-4 accuracy of DMS-Deep exceed the Top-10 accuracy of all search methods that utilize iterative application of single-step models, highlighting the advantages of the multistep-first approach. To understand this performance difference, it helps to notice that iterative application of single-step calls results in a breadth-first search-like exploration of a node, as all reactants leading to a compound are generated at the same time. Our approach, however, is a depth-first search-like: the whole left sub-tree of target compound in Fig. 1 is predicted before the immediate right child. This provides an extra context for the prediction of the second precursor to the target compound (or any other right child), which reduces the search space and finds more relevant compounds faster.
A few words on the relevance of Top-$K$ accuracy as the primary evaluation metric are in order. One might question whether the number of targets for which valid (i.e., having valid SMILES) routes are found (or solvability rate if all SMs are in a given stock set) and the average length of routes predicted should be considered. It is essential to realize that even if you ensure that every single-step transformation corresponds to a known reaction by template matching, the route may still be experimentally not feasible simply because atoms that are unchanged (i.e., the ones ignored by the template) contribute to the energies of the reacting orbitals through inductive and resonance effects, thereby changing the energetic favorability of the reaction. As a result, short of performing highly-accurate ab initio calculations of energies (after corresponding geometry optimization) with all relevant solvation effects and considering thermal corrections, there is no computational metric that would assess experimental feasibility of the predicted route. In addition, one could argue that the ability to discover an experimentally verified route is a prerequisite for finding shorter, but still feasible, routes. Therefore, Top-$K$ accuracy is the closest proxy to route quality. Additionally, the ability to predict correct routes with better ranks is a significant improvement because it reduces the number of candidates to be passed through the aforementioned highly expensive energy calculations. Finally, one could argue that single-step methods do not provide inherent ranking of candidates, so the performance of search methods may depend on the choice of ranking metric. However, that is yet another benefit of our approach: there is no need in devising another algorithm or model for route ranking, the beam probabilities are sufficiently accurate.
The superior performance of DMS models does not incur additional computational cost. A search with a beam size of 50 takes 3-16 seconds (depending on the number of steps) with a DMS-Flash or a DMS-Flex model on a single GPU (NVIDIA M4000), which is comparable to the time required to find the first successful route using previous methods (7-50 seconds [44]).
@@
Figure 3: Distribution of Top-1 and Top-10 accuracy of predictions with DMS-Flex (Duo) on test sets n1 and n5. There is only one route with length 10 in n1, and DMS-Flex (Duo) does not predict it correctly. That route is reproduced by splitting it in half, as shown in Fig. 4
3.1.2 MoE Transformers
The DMS-Flex (Flexible), a mixture-of-experts model trained with two out of three experts active, provides the largest performance improvement when evaluated with two experts active (Duo mode): 2.5x increase in Top-1 on n1 and 3.9x on n5. Evaluating DMS-Flex with one expert active (Mono) mode worsens the performance by 30%; given that the number of valid routes (Tables S1-S2) does not change significantly, one could consider the Mono mode as a source of diversity in predictions. Notably, even though the DMS-Flex is trained with SM, it can be evaluated without providing SM (empty padding tokens are fed into the encoder instead) which still results in 34% Top-1 accuracy and 47+Top-10. Evaluating DMS-Flex without providing the number of steps as an input drops performance significantly, but a different model with the same architecture trained without number of steps as an input, DMS-Explorer (Fig. S1 shows the distribution of route lengths predicted by Explorer relative to the length of an experimental route), results in only a 8% drop in accuracy.
Using a larger DMS-Wide model trained without SM provides comparable performance to that of DMS-Deep. An important advantage of DMS-Wide is that tokens passed through different active experts can be processed on a GPU in parallel, which enables faster predictions. In addition, DMS-Wide has fewer layers, so fewer consecutive calculations have to be performed; in practice, doubling the number of encoder layers results in doubling of generation time. One might also conclude that similar performance of DMS-Deep and DMS-Wide, which have comparable parameter counts, implies that the expressive power mostly depends on the number of parameters, and not on the extra non-linearities provided by larger number of layers.
Our experimentation with the increase in the number of experts, the number of encoder/decoder layers, or the feed-forward multiplier evaluated on small subsets of n1 or n5, show no significant improvement in performance (Tables S4-S7), which most likely means that the 20 M (with SM) or 40 M (without SM) are optimal given the current size of the training set.
Fig. 3 shows the distribution of Top-1 and Top-10 accuracy over different route lengths on both n1 and n5. Given that 85% of routes in the training partition have 4 or fewer steps Fig. 2), one would expect the accuracy to decrease dramatically with increasing route length. However, the performance on routes with 5-8 steps is comparable to that of shorter routes, and the performance on 9-step routes is even comparable to that of 2-step routes. The exceptional accuracy on 9-step routes could be rationalized by the fact that their relative abundance in n1 and n5 does not exceed the relative abundance in the training partition (unlike for routes with length 3-7, as seen in Fig. 3). Top-1 and Top-10 accuracies of other DMS models for different route lengths are given in Tables S8-S11.